OpenStudy (anonymous):

What is the slope of the line passing through the points (1, –5) and (4, 1)? A.2 B.-4/5 C.5/4 D.-2

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

will fan and medal

OpenStudy (ariane12):

2

OpenStudy (anonymous):

thank you so much

OpenStudy (ariane12):

M=y2-y1/x2-x1 =1-(-5)/4-1 =6/3 =2